# Dimension-independent convergence rates of randomized nets using median-of-means

### v2

@@ 2.4 Walsh coefficients and function smoothness
  • $\lceil\kappa\rceil_{q}$: The $q$-th largest element of $\kappa$ (or 00 if $|\kappa|<q$).
@@
[…]
@@
$W_{\kappa}(x)=\int_{[0,1]}(-1)^{\vec{x}(\lfloor\kappa\rfloor)}W_{\kappa\setminus\lfloor\kappa\rfloor}(x)\,\mathrm{d}x.$
@@
Equation (6) comes from [30, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [30, Section 3]. ∎
For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $[0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$. In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3. This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (6), as discussed in [25, Remark 1].
@@ 3 ANOVA decomposition and effective dimensionality
To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [19]:
@@
To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [3]. For a function $f:[0,1]^{d}\to\mathbb{R}$ and a subcube $J=\prod_{j=1}^{d}[a_{j},b_{j})$, we define the alternating sum over $J$:
@@
where the supremum is over all partitions $\mathcal{P}$ of $[0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$. When $d=0$ and $f\in\mathbb{R}$, $V^{(0)}_{\lambda}(f)=|f|$ by convention.
The Vitali variation generalizes Sobolev norms: for $f$ with weak mixed derivative $f^{(1,\dots,1)}\in L^{p}([0,1]^{d})$, [20] shows $V_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$. Crucially, $V^{(d)}_{\lambda}(f)$ can be finite even for discontinuous $f$ (see [25, Section 4.1] for an example).
@@
By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [25, Section 3.2],
@@
implying $V^{v}_{\lambda}(f)<\infty$ if almost every section $f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})$ has bounded variation.
@@
Let $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})<\infty$ for nonempty $u\subseteq 1{:}s$, $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and a subset $v\subseteq\{j\mid\alpha_{j}=\alpha\}$. Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$,
@@
$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=$
where
From the proof of [25, Theorem 3], we can map $v$ to $1{:}|v|$ and show
and $\lceil\kappa_{j}\rceil_{\alpha+1}=0$ for $j\in u\setminus v$. ∎
@@
where $f_{u}$ is the ANOVA component of $f$ defined in equation (8). Notice that we add the restriction $\alpha_{j}>0$ for $j\in u\setminus v$ because $k_{j}\in\mathbb{N}$ implies $\alpha_{j}=|\kappa_{j}|>0$ in the statement of Lemma 4. In particular, $\|f\|_{u,\alpha,\lambda}=V^{u}_{\lambda}(f_{u})$ when $\alpha=0$. We further set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
@@
$\gamma_{u}=\frac{\|f\|_{u,\alpha,\lambda}}{\|f\|_{\alpha,\lambda}},$ (16)
which satisfies $\sup_{u\subseteq 1{:}s}\gamma_{u}=1$ and $\gamma_{\varnothing}=0$. The decay of $\gamma_{u}$ encodes effective dimensionality, with $\gamma_{u}=1$ for all nonempty $u\subseteq 1{:}s$ representing the worst dimensionality. As shown in Theorem 4, dimension-independent convergence rates emerge under suitable decay conditions on $\gamma_{u}$.
@@ 4 Tractability under the completely random design
First, we define $\|\kappa\|=\sum_{\ell\in\kappa}\ell$ for $\kappa\subseteq\mathbb{N}$, with $\|\varnothing\|=0$. Motivated by Lemma 4, we choose $K$ in equation (5) for given $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$ to be:
@@
we can apply Lemma 4 and equation (15) to get
we have for any $\theta\in(0,1)$ and $\rho_{\theta}=4^{-1+\theta}$
$g(t)=\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha^{\prime}},$
and let $[t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [6] for background knowledge on generating functions). It follows that $[t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$. Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
Our conclusion follows once we put the above bounds into equation (19). ∎
@@
with the sum from $1$ to $\alpha$ interpreted as 00 when $\alpha=0$.
@@
Next, we notice
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
Let $\overline{N}=\lfloor(T-N)/(\alpha+\lambda)\rfloor$. For nonempty $v\subseteq u$ and integer $n\geqslant|v|$, there are (${n-1\choose|v|-1}$ number of $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfying $\|\boldsymbol{\ell}_{v}\|_{1}=n$. Hence, when $\overline{N}\geqslant|v|$,
$\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}=$ $\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}$ (24)
$=$ $\sum_{N=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n+1}-\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}$
$=$ (${\overline{N}\choose|v|-1}2^{\overline{N}+1}-\sum_{n=|v|}^{\overline{N}}\Big{\{}{n-1\choose|v|-1}-{n-2\choose|v|-1}\Big{\}}2^{n}$
$\leqslant$ (${\overline{N}\choose|v|-1}2^{\overline{N}+1}.$
When $\overline{N}<|v|$, no $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfies $\|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)$ and the above sum is trivially 00. Putting the above bound into equation (23), we get
To bound the sum over $N$, consider the generating function
$g(t)=\prod_{j\in u}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}.$
for $t=2^{-1/(\alpha+\lambda)}$ and
$=$ $A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|},$
$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}.$
Putting the above bounds into equation (25), we finally get
$|K_{u}(T)|\leqslant$ $2^{T/(\alpha+\lambda)}\Big{(}B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}A_{\alpha,\lambda}^{|v|}B^{|u|-|v|}_{\alpha,\lambda}\Big{)}$ (27)
$=$ $2^{T/(\alpha+\lambda)}\Big{(}B_{\alpha,\lambda}+\max\big{(}T/(\alpha+\lambda),1\big{)}A_{\alpha,\lambda}\Big{)}^{|u|}.\qed$ (28)
@@
Let $\delta=1/16$ in Lemma 2 and Corollary 1. By Lemma 2, equation (5) and $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for $\boldsymbol{k}\neq\boldsymbol{0}$, it suffices to show
@@
Given a failure probability tolerance $\varepsilon>0$, we can set $r=\lceil\log_{2}(\varepsilon^{-1})\rceil$ to ensure $2^{-r}\leqslant\varepsilon$. For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, where $N=r2^{m}$ is the total number of function evaluations.
To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$, so $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$. Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}\leqslant 1$,
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big{)}\leqslant 2^{-r}\leqslant\varepsilon,$ (33)
where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is almost optimal [25, Subsection 3.2].
@@
In settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$, dimension-independent error bounds hold if there exist weights $\{\gamma_{j},j\in 1{:}s\}$ satisfying
$\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}\quad\forall u\subseteq 1{:}s\quad\text{ and }\quad\sum_{j=1}^{s}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\leqslant C_{\infty}$ (34)
for a constant $C_{\infty}$ independent of $s$. To see this, we first bound
@@
which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [11, Lemma 3]. Putting the above bounds into equation (31), we conclude that equation (33) holds with a constant $C_{\eta}$ independent of $s$. Notably, equation (34) resembles the tractability conditions assumed by [9, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
@@ 5 Tractability under more general randomization
While the completely random design offers theoretical advantages in near-optimal convergence rates and tractability, the random linear scrambling often outperforms it in practical settings due to tighter error bounds on low-dimensional ANOVA components. For example, [12] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets. In this section, we extend the analysis of Section 4 to a broader class of randomization schemes, with the random linear scrambling as a key example.
@@
From [9, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [9, Section 2]. We frame this assumption to align with broader applications in higher-order digital nets [4], where the parameter $d$ (the order of digital nets) may exceed $1$. This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
with the sum from $1$ to $\alpha-d$ interpreted as 00 when $\alpha=d$.
@@
We adopt the notation used in the proof of Lemma 6. For $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$, we again consider the one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}$ for $j\in v$. Because $d\in\mathbb{N}$ and $d<\alpha+\lambda$, we must have $d\leqslant\alpha$. By $|\kappa^{+}_{j}|=\alpha\geqslant d$ for $j\in v$,
and
Let
Similar to equation (23), we can write
$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$ (36)
$N_{*}=\frac{(\alpha+\lambda)T^{\prime}-dT}{\alpha+\lambda-d}.$
Therefore,
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1},$
Next, we notice
the number of possible $(\kappa^{+}_{j,d},j\in w)$ for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$ with a given set of $(\kappa^{-}_{j,d},j\in u)\in\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$ is bounded by the number of $(\kappa^{+}_{j,d},j\in w)$ with $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=N-N^{\prime}$, which is $|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ by definition. This implies $|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\times|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ and
$=$ $D_{\alpha,\lambda,d,\theta^{\prime}}^{|u|},$
where $D_{\alpha,\lambda,d,\theta^{\prime}}$ is defined in equation (35). ∎
@@
$\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$ (39)
@@
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}2^{m-\tilde{t}_{u}}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
@@
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
@@
Our conclusion follows from Lemma 2 and equation (5) after plugging in $T_{u,m}$ from equation (40). ∎
@@
Consider again settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$. We assume equation (34) holds for weights $\{\gamma_{j},j\in 1{:}s\}$. Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C_{\eta}$ independent of $s$. Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [31] that the base-2 Niederreiter sequence [22] satisfies
$t_{u}\leqslant\sum_{j\in u}\log_{2}(j)+\log_{2}\log_{2}(j+2)+2,$ (42)
while the Sobol’ sequence [29] satisfies
$t_{u}\leqslant\sum_{j\in u}\log_{2}(j)+\log_{2}\log_{2}(j+1)+\log_{2}\log_{2}\log_{2}(j+3)+C$
@@
For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for nonnegative numbers $\{\tilde{t}_{j},j\in 1{:}s\}$. We can then bound
@@
$\lim_{s\to\infty}\sum_{j=1}^{s}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$

### v3

@@ 2.4 Walsh coefficients and function smoothness
  • $\lceil\kappa\rceil_{q}$: The $q$-th largest element of $\kappa$ (or 0 if $|\kappa|<q$).
@@
The following lemma connects the Walsh coefficients $\hat{f}(\boldsymbol{k})$ to the smoothness of $f$. For $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{s})\in\mathbb{N}^{s}_{0}$, we define the mixed partial derivative:
@@
$W_{\kappa}(x)=\int_{[0,1]}(-1)^{\vec{x}(\lfloor\kappa\rfloor)}W_{\kappa\setminus\lfloor\kappa\rfloor}(x)\,\mathrm{d}x$
@@
Equation (5) comes from [SUZUKI20161, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [SUZUKI20161, Section 3]. ∎
For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$. In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3. This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (5), as discussed in [pan2024automatic, Remark 1].
@@ 3 ANOVA decomposition and effective dimensionality
To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [meandim]:
@@
To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [dick:2008], which generalizes the usual Sobolev norms. For a function $f:[0,1]^{d}\to\mathbb{R}$ and a subcube $J=\prod_{j=1}^{d}[a_{j},b_{j})$, we define the alternating sum over $J$:
@@
where the supremum is over all partitions $\mathcal{P}$ of $0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$. When $d=0$ and $f\in\mathbb{R}$, $V^{(0)}_{\lambda}(f)=|f|$ by convention.
For $f$ with weak mixed derivative $f^{(1,\dots,1)}$, we have the inequality
$V^{(d)}_{1}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{2}},$ (10)
with equality attained when $f\in C^{(1,\dots,1)}([0,1]^{s})$ [dick:2008, equation (3.4)]. More generally, [LIU2025101935] shows $V^{(d)}_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$. Discontinuous $f$ can also have finite $V^{(d)}_{\lambda}(f)$ (see [pan2024automatic, Section 4.1] for an example).
@@
By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [pan2024automatic, Section 3.2],
@@
which, together with equation (10), implies
$V^{v}_{1}(f)\leqslant\|f^{(\boldsymbol{1}_{v},\boldsymbol{0}_{v^{c}})}\|_{L^{2}}$ (13)
if the weak derivative $f^{(\boldsymbol{1}_{v},\boldsymbol{0}_{v^{c}})}$ exists. Here $(\boldsymbol{1}_{v},\boldsymbol{0}_{v^{c}})$ denotes the vector whose $j$-th entry equals $1$ if $j\in v$ and equals 0 otherwise.
@@
Let $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})$$\infty$ for nonempty $u\subseteq 1{:}s$, $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and a subset $v\subseteq\{j\mid\alpha_{j}=\alpha\}$. Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$,
@@
See Appendix A. ∎
@@
where $f_{u}$ is the ANOVA component of $f$ defined in equation (7). Notice that we add the restriction $\alpha_{j}>0$ for $j\in u\setminus v$ because $k_{j}\in\mathbb{N}$ implies $\alpha_{j}=|\kappa_{j}|>0$ in the statement of Lemma 4. In particular, $\|f\|_{u,0,\lambda}=V^{u}_{\lambda}(f_{u})$. When $\lambda=1$ and the weak mixed derivative $f^{(\alpha+1,\dots,\alpha+1)}$ exists, equation (13) implies the bound
$\|f\|_{u,\alpha,1}\leqslant\sup_{\boldsymbol{\alpha}_{u}\in\{1:(\alpha+1)\}^{|u|}}\|f_{u}^{(\boldsymbol{\alpha}_{u})}\|_{L^{2}}.$
We further set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
@@
$\gamma_{u}=\gamma_{u}(f,\alpha,\lambda)=\frac{\|f\|_{u,\alpha,\lambda}}{\|f\|_{\alpha,\lambda}},$ (16)
which satisfies $\sup_{u\subseteq 1{:}s}\gamma_{u}=1$ and $\gamma_{\varnothing}=0$. The relative variation $\gamma_{u}$ depends on $f,\alpha$ and $\lambda$, but we suppress this dependence in the notation as these variables are fixed throughout our theoretical analysis. Section 6 later exemplifies how $\gamma_{u}$ changes under different choices of $\alpha$.
The decay of $\gamma_{u}$ encodes effective dimensionality, with $\gamma_{u}=1$ for all nonempty $u\subseteq 1{:}s$ representing the worst dimensionality. As shown in Theorem 5, dimension-independent convergence rates emerge under suitable decay conditions on $\gamma_{u}$.
@@ 4 Tractability under the completely random design
First, we define $\|\kappa\|=\sum_{\ell\in\kappa}\ell$ for $\kappa\subseteq\mathbb{N}$, with $\|\varnothing\|=0$. Motivated by Lemma 4, we choose $K$ in Lemma 2 for given $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$ to be:
@@
See Appendix B. ∎
@@
with the sum from 11 to $\alpha$ interpreted as 0 when $\alpha=0$.
@@
See Appendix C. ∎
@@
Let $\delta=1/16$ in Lemma 2 and Corollary 1. By Lemma 2 and $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for $\boldsymbol{k}\neq\boldsymbol{0}$, it suffices to show
@@
Given a failure probability tolerance $\bar{\delta}>0$, we can set $r=\lceil\log_{2}(\bar{\delta}^{-1})\rceil$ to ensure $2^{-r}\leqslant\bar{\delta}$. For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, where $N=r2^{m}$ is the total number of function evaluations.
To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$ because $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$. Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}<\infty$ that
$\Pr\Big(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}\|f\|^{2}_{\alpha,\lambda}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big)\leqslant 2^{-r}\leqslant\bar{\delta},$ (24)
where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is nearly optimal, as it aligns with the lower bound established in [bakh:1959] for one-dimensional $f$ with $\lambda$-Hölder continuous $f^{(\alpha)}$, a subclass of functions with finite $\|f\|_{\alpha,\lambda}$.
@@
In settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$, dimension-independent error bounds hold if there exist weights $\{\gamma_{j},j\in\mathbb{N}\}$ satisfying
$\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}\quad\forall u\subseteq\mathbb{N}\quad\text{ and }\quad C_{\infty}=\sum_{j=1}^{\infty}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}<\infty.$ (25)
To see this, we first bound
@@
which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [HICKERNELL2003286, Lemma 3]. Putting the above bounds into equation (22), we conclude that equation (24) holds with a constant $C_{\eta}$ independent of $s$. Notably, equation (25) resembles the tractability conditions assumed by [Goda2024, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
@@ 5 Tractability under more general randomization
While the completely random design offers theoretical advantages in near-optimal convergence rates and tractability, the random linear scrambling often outperforms it in practical settings due to tighter error bounds on low-dimensional ANOVA components. For example, [joe:kuo:2008] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets. In this section, we extend the analysis of Section 4 to a broader class of randomization schemes, with the random linear scrambling as a key example.
@@
From [Goda2024, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [Goda2024, Section 2]. We frame this assumption to align with broader applications in higher-order digital nets [dick:2011], where the parameter $d$ (the order of digital nets) may exceed $1$. This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
with the sum from 11 to $\alpha-d$ interpreted as 0 when $\alpha=d$.
@@
See Appendix D. ∎
@@
[…]
@@
[…]
@@
[…]
@@
Our conclusion follows from Lemma 2 after plugging in $T_{u,m}$ from equation (28). ∎
@@
Consider again settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$. We assume equation (25) holds for weights $\{\gamma_{j},j\in\mathbb{N}\}$. Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C_{\eta}$ independent of $s$. Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [c812a4a5-75ac-3614-936d-5782e44941d0] that the base-2 Niederreiter sequence [NIEDERREITER198851] satisfies
$t_{u}\leqslant\sum_{j\in u}\Big(\log_{2}(j)+\log_{2}\log_{2}(j+2)+2\Big),$ (30)
while the Sobol’ sequence [sobol67] satisfies
$t_{u}\leqslant\sum_{j\in u}\Big(\log_{2}(j)+\log_{2}\log_{2}(j+1)+\log_{2}\log_{2}\log_{2}(j+3)+C\Big)$
@@
For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for nonnegative numbers $\{\tilde{t}_{j},j\in\mathbb{N}\}$. We can then bound
@@
$\sum_{j=1}^{\infty}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$